# Elastic Softening in Synthetic Diamonds

### v2

@@ Abstract
In summary, our ultrasound studies clearly reveal the presence of neutral single vacancies on the ppb level in all three diamond types investigated. In particular, we have uncovered that CVD diamonds, which are widely used for optical measurements in quantum technologies, also have atomic vacancies on the ppb level comparable to HPHT-grown diamonds. This result should be considered for future downsizing and removal of spin-noise sources in NV centers used as quantum devices. Importantly, the present study focused on atomic vacancies formed spontaneously during single-crystal growth, excluding those formed by neutron- or electron-beam irradiation. These results are of particular importance for researchers utilizing such diamonds in various fields, such as in quantum information, biological sensing, and power devices. Furthermore, our findings provide a novel perspective for studying atomic vacancies in diamonds. Further investigations of the elastic response of various types of diamonds at cryogenic temperature are required to explore in detail the underlying physics of diamond vacancies with respect to possible technological applications.
@@ Acknowledgements.
- Kasu et al. [2021] M. Kasu, N. C. Saha, T. Oishi, and S.-W. Kim, Fabrication of diamond modulation-doped FETs by NO delta doping in an AlO gate layer, Appl. Phys. Express 14, 051004 (2021).
@@
- Sumiya et al. [2015] H. Sumiya, K. Harano, and K. Tamasaku, HPHT synthesis and crystalline quality of large high-quality (001) and (111) diamond crystals, Diamond and Related Materials 58, 221 (2015).
@@
- Luo et al. [2022] T. Luo, L. Lindner, J. Langer, V. Cimalla, X. Vidal, F. Hahl, C. Schreyvogel, S. Onoda, S. Ishii, T. Ohshima, D. Wang, D. A. Simpson, B. C. Johnson, M. Capelli, R. Blinder, and J. Jeske, Creation of nitrogen-vacancy centers in chemical vapor deposition diamond for sensing applications, New J. Phys. 24, 033030 (2022).
- [20] See Supplemental Material for (1) Thermodynamic Consideration of the Formation of Atomic Vacancies, (2) Neutral Single Vacancy V in Silicon and Diamond, (3) Experimental Details, (4) Ultrasonic Method, (5) Wave Functions for Hybrid Orbitals and Active Electric Quadrupole, (6) Quadrupolar Susceptibility, (7) Comparison of Si and diamond, (8) Magnetic Field Dependence of the Softening of , (9) Inconsistency of Quantum Ground State between Non-Irradiated and irradiated Diamond, and (10) Effect of NV, NV, P Centers, and Nuclear Spin.
- Bogdanov et al. [2018] K. V. Bogdanov, M. V. Zhukovskaya, V. Y. Osipov, E. V. Ushakova, M. A. Baranov, K. Takai, A. Rampersaud, and A. V. Baranov, Highly intensive emission of the NV centers in synthetic HPHT microdiamonds at low nitrogen doping, APL Materials 6, 086104 (2018).
- Subedi et al. [2021] S. Subedi, V. Fedorov, S. Mirov, and M. Markham, Spectroscopy of GR1 centers in synthetic diamonds, Optical Materials Express 11, 757 (2021).
- Ref [2022] Handbook of Scientific Tables (World Scientific, 2022).
@@
- Yamada-Kaneta et al. [2006] H. Yamada-Kaneta, T. Goto, Y. Saito, Y. Nemoto, K. Sato, K. Kakimoto, and S. Nakamura, Vacancies in defect-free zone of point-defect-controlled CZ silicon observed by low-temperature ultrasonic measurements, Materials Science and Engineering: B 134, 240 (2006).
@@
- Goto et al. [2007] T. Goto, H. Yamada-Kaneta, K. Sato, M. Hikin, Y. Nemoto, and S. Nakamura, Observation of vacancy in crystalline silicon using low-temperature ultrasonic measurements, Physica B: Condensed Matter 401-402, 109 (2007).
- McSkimin and Andreatch [1972] H. J. McSkimin and P. Andreatch, Elastic Moduli of Diamond as a Function of Pressure and Temperature, Journal of Applied Physics 43, 2944 (1972).
- Migliori et al. [2008] A. Migliori, H. Ledbetter, R. G. Leisure, C. Pantea, and J. B. Betts, Diamond’s elastic stiffnesses from 322 K to 10 K, Journal of Applied Physics 104, 053512 (2008).
- Nagakubo et al. [2016] A. Nagakubo, M. Arita, H. Ogi, H. Sumiya, N. Nakamura, and M. Hirao, Elastic constant of C diamond between 10 and 613 K, Appl. Phys. Lett. 108, 221902 (2016).
- Laszlo et al. [2010] I. Laszlo, M. Kertesz, B. Slepetz, and Y. Gogotsi, Simulations of large multi-atom vacancies in diamond, Diamond and Related Materials 19, 1153 (2010).
- Kamihara et al. [2022] D. Kamihara, T. Shimizu, and K. Uchida, Discovery of Peculiar Electronic Structures of Decavacancy V in Silicon Crystal, J. Phys. Soc. Jpn. 91, 064709 (2022).
- Cox et al. [1994] A. Cox, M. E. Newton, and J. M. Baker, C,N and N ENDOR measurements on the single substitutional nitrogen centre (P1) in diamond, Journal of Physics: Condensed Matter 6, 551 (1994).
@@
- Davies [1979] G. Davies, Dynamic jahn-teller distortions at trigonal optical centres in diamond, Journal of Physics C: Solid State Physics 12, 2551 (1979).
- Prentice [2017] J. C. A. Prentice, First-principles study of the dynamic jahn-teller distortion of the neutral vacancy in diamond, Physical Review B 95, 014108 (2017).
- Braukmann [2018] D. Braukmann, Circularly polarized zero-phonon transitions of vacancies in diamond at high magnetic fields, Physical Review B 97, 195448 (2018).
- Hayami [2018] S. Hayami, Classification of atomic-scale multipoles under crystallographic point groups and application to linear response tensors, Phys. Rev. B 98, 165110 (2018).
- Doherty et al. [2013] M. W. Doherty, N. B. Manson, P. Delaney, F. Jelezko, J. Wrachtrup, and L. C. Hollenberg, The nitrogen-vacancy colour centre in diamond, Physics Reports 528, 1 (2013).
- Schlüter [1985] M. Schlüter, Proc. Int. School of Physics “Enrico Fermi”, Proc. Int. School of Physics , p. 495 (1985).

### v4

@@ V Conclusion
In summary, our ultrasound studies strongly suggest the presence of the positive vacancies V+. However, we cannot rule out the existence of other unresolved vacancies or interstitial states, which have electric quadrupolar degrees of freedom on the sub-ppb level in all three diamond types investigated. In particular, we have uncovered that CVD diamonds, which are widely used for optical measurements in quantum technologies, also have such V+ vacancies with concentration comparable to HPHT-grown diamonds. This result should be considered for future downsizing and removal of spin noise sources in NV centers used as quantum devices. Importantly, the present study focused on the V+ sites formed spontaneously during single-crystal growth, excluding those formed by neutron or electron-beam irradiation damages. These results are of particular importance for researchers utilizing such diamonds in various fields, such as in quantum information, biological sensing, and power devices. Furthermore, our findings provide a novel perspective for studying unknown quantum ground states of the vacancies in diamonds. Further investigations of the elastic response of various types of diamonds at cryogenic temperature are required to explore in detail the underlying physics of diamond vacancies or defects with respect to possible technological applications.
@@ Acknowledgements.
- Kasu et al. [2021] M. Kasu, N. C. Saha, T. Oishi, and S.-W. Kim, Fabrication of diamond modulation-doped FETs by NO2 delta doping in an Al2O3 gate layer, Appl. Phys. Exp. 14, 051004 (2021).
@@
- Sumiya et al. [2015] H. Sumiya, K. Harano, and K. Tamasaku, HPHT synthesis and crystalline quality of large high-quality (001) and (111) diamond crystals, Dia. Rel. Mater. 58, 221 (2015).
@@
- [16] See Supplemental Material for (1) Thermodynamic Consideration of the Formation of Atomic Vacancies, (2) Single Vacancies V0 and V+ in Silicon and Diamond, (3) Ultrasonic Method, (4) Wave Functions for sp3 Hybrid Orbitals and Active Electric Quadrupole, (5) Quadrupolar Susceptibility, (6) Comparison of Si and Diamond, and (7) Magnetic Field Dependence of the Softening of .
- Bogdanov et al. [2018] K. V. Bogdanov, M. V. Zhukovskaya, V. Y. Osipov, E. V. Ushakova, M. A. Baranov, K. Takai, A. Rampersaud, and A. V. Baranov, Highly intensive emission of the NV- centers in synthetic HPHT microdiamonds at low nitrogen doping, APL Materials 6, 086104 (2018).
- Subedi et al. [2021] S. Subedi, V. Fedorov, S. Mirov, and M. Markham, Spectroscopy of GR1 centers in synthetic diamonds, Opt. Mater. Exp. 11, 757 (2021).
- Baldwin [1963] J. A. Baldwin, Electron Paramagnetic Resonance Investigation of the Vacancy in Diamond, Phys. Rev. Lett. 10, 220 (1963).
- Loubser and Wyk [1978] J. H. N. Loubser and J. A. v. Wyk, Electron Spin Resonance in the Study of Diamond, Rep. Prog. Phys. 41, 1201 (1978).
- Mainwood and Stoneham [1997] A. Mainwood and A. M. Stoneham, Stability of electronic states of the vacancy in diamond, J. Phys.: Cond. Matt. 9, 2453 (1997).
- Isoya [1992] J. Isoya, EPR identification of the negatively charged vacancy in diamond, Phys. Rev. B 45, 1436 (1992).
@@
- Luo et al. [2022] T. Luo, L. Lindner, J. Langer, V. Cimalla, X. Vidal, F. Hahl, C. Schreyvogel, S. Onoda, S. Ishii, T. Ohshima, D. Wang, D. A. Simpson, B. C. Johnson, M. Capelli, R. Blinder, and J. Jeske, Creation of nitrogen-vacancy centers in chemical vapor deposition diamond for sensing applications, New J. Phys. 27, 033030 (2022).
- Rik [2022] Handbook of Scientific Tables (World Scientific, 2022).
@@
- Davies [1981] G. Davies, The Jahn-Teller effect and vibronic coupling at deep levels in diamond, Rep. Prog. Phys. 44, 787 (1981).
- McSkimin and Andreatch [1972] H. J. McSkimin and P. Andreatch, Elastic Moduli of Diamond as a Function of Pressure and Temperature, J. Appl. Phys. 43, 2944 (1972).
- Migliori et al. [2008] A. Migliori, H. Ledbetter, R. G. Leisure, C. Pantea, and J. B. Betts, Diamond’s elastic stiffnesses from 322 K to 10 K, J. Appl. Phys. 104, 053512 (2008).
- Nagakubo et al. [2016] A. Nagakubo, M. Arita, H. Ogi, H. Sumiya, N. Nakamura, and M. Hirao, Elastic constant of 12C diamond between 10 and 613 K, Appl. Phys. Lett. 108, 221902 (2016).
- Yamada-Kaneta et al. [2006] H. Yamada-Kaneta, T. Goto, Y. Saito, Y. Nemoto, K. Sato, K. Kakimoto, and S. Nakamura, Vacancies in defect-free zone of point-defect-controlled CZ silicon observed by low-temperature ultrasonic measurements, Mater. Sci. and Eng.: B 134, 240 (2006).
@@
- Goto et al. [2007] T. Goto, H. Yamada-Kaneta, K. Sato, M. Hikin, Y. Nemoto, and S. Nakamura, Observation of vacancy in crystalline silicon using low-temperature ultrasonic measurements, Physica B 401-402, 109 (2007).
- Davies [1994] G. Davies, ed., Properties and growth of diamond, EMIS datareviews series 9 (INSPEC, 1994).
@@
- Prentice [2017] J. C. A. Prentice, First-principles study of the dynamic Jahn-Teller distortion of the neutral vacancy in diamond, Phys. Rev. B 95, 014108 (2017).
@@
- Davies [1979] G. Davies, Dynamic Jahn-Teller distortions at trigonal optical centres in diamond, J. Phys. C: Solid State Phys. 12, 2551 (1979).
- Davies [1992] G. Davies, Vacancy-related centers in diamond, Phys. Rev. B 46, 13157 (1992).
- Coulson and Larkins [1971] C. Coulson and F. Larkins, Isolated single vacancy in diamond—I. Electronic structure, J. Phys. Chem. Solids 32, 2245 (1971).
- Siyushev [2013] P. Siyushev, Optically Controlled Switching of the Charge State of a Single Nitrogen-Vacancy Center in Diamond at Cryogenic Temperatures, Phys. Rev. Lett. 110, 167402 (2013).
- Meara [2019] C. J. Meara, Density functional theory study on magnetically detecting positively charged nitrogen-vacancy center in diamond, Phys. Rev. B 100, 104108 (2019).
- Goto [2004] T. Goto, Tunneling and rattling in clathrate crystal, Phys. Rev. B 70, 184126 (2004).
- Araki et al. [2012] K. Araki, T. Goto, K. Mitsumoto, Y. Nemoto, M. Akatsu, H. S. Suzuki, H. Tanida, S. Takagi, S. Yasin, S. Zherlitsyn, and J. Wosnitza, Dissipation in Non-Kramers Doublet of PrMg3, J. Phys. Soc. Jpn. 81, 023710 (2012).
- Kamihara et al. [2022] D. Kamihara, T. Shimizu, and K. Uchida, Discovery of Peculiar Electronic Structures of Decavacancy V10 in Silicon Crystal, J. Phys. Soc. Jpn. 91, 064709 (2022).
- Cox et al. [1994] A. Cox, M. E. Newton, and J. M. Baker, 13C,14N and 15N ENDOR measurements on the single substitutional nitrogen centre (P1) in diamond, J. Phys.: Cond. Matt. 6, 551 (1994).
- Doherty et al. [2013] M. W. Doherty, N. B. Manson, P. Delaney, F. Jelezko, J. Wrachtrup, and L. C. Hollenberg, The nitrogen-vacancy colour centre in diamond, Phys. Rep. 528, 1 (2013).
@@
- van Wyk [1995] J. A. van Wyk, Magnetic-resonance measurements on the excited state of the neutral vacancy in diamond, Phys. Rev. B 52, 12657 (1995).